# An Evidence-Based Post-Hoc Adjustment Framework for Anomaly Detection Under Data Contamination

### v1

@@ B.3.1 Evidence for visual datasets
For the evidence function in image-based AD, we propose using Contrastive Language-Image Pre-training (CLIP) (Radford et al.,, 2021), a robust large-scale framework that learns joint vision-language representations from web-collected image-text pairs. While CLIP has been explored in prior work as a zero-shot AD method (Jeong et al.,, 2023; Zhou et al.,, 2024), its performance varies across different datasets. Although CLIP excels in detecting anomalies in real-world image datasets such as CIFAR10, it faces significant challenges when applied to domain-specific datasets, particularly those used for industrial inspection, like MVTec. This limitation stems from the lack of domain-specific knowledge in CLIP’s pre-training. In this section, we describe how CLIP is integrated into EPHAD as an evidence function $T(x)$, leveraging its strengths while mitigating its limitations in specialized domains.
Given a dataset $\mathcal{D}:=\{(x_{j},t_{j})\}_{j=1}^{n}$, CLIP trains an image encoder $e_{i}$ and a text encoder $e_{t}$ using contrastive learning (Chen et al.,, 2020), maximizing the cosine similarity between $e_{i}(x_{j})$ and $e_{t}(t_{j})$ for all $(x_{j},t_{j})\in\mathcal{D}$. For an input image $x$, CLIP performs zero-shot classification (Radford et al.,, 2021) by computing a $k$-way categorical distribution over a set of candidate class texts ${\mathcal{C}}=\{c_{1},\dots,c_{k}\}$
@@
Building on Jeong et al., (2023), we use CLIP as evidence function $T(x)$ in EPHAD. We start by defining two lists of textual prompt templates, ${\mathcal{T}}_{N}=\{n_{1},\cdots,n_{k}\}$ and ${\mathcal{T}}_{A}=\{a_{1},\cdots,a_{k}\}$, corresponding to normal and anomalous classes, respectively. The list of prompts is provided in Table 3. These templates are dataset-dependent, reflecting subjectivity (e.g., “missing wire" as anomalous for cables). For each label, we generate two lists of prompts for normal and anomalous cases using ${\mathcal{T}}_{N}$ and ${\mathcal{T}}_{A}$ and compute the mean of text embeddings $t_{N}$ and $t_{A}$. Finally, given an input image $x$, the evidence $T(x)$ during test-time is computed as:
$T(x):=\frac{\exp\left(\langle e_{i}(x),t_{A}\rangle/\gamma\right)}{\exp\left(\langle e_{i}(x),t_{N}\rangle/\gamma\right)+\exp\left(\langle e_{i}(x),t_{A}\rangle/\gamma\right)}.$
One potential concern when using pre-trained models like CLIP is the overlap between their training data and the test samples encountered in downstream tasks. Such overlap could challenge the assumption that test-time statistics are based solely on test data. However, Radford et al., (2021) provides an extensive analysis of this issue and shows that excluding all overlapping samples from CLIP’s pre-training corpus leads to only a negligible performance drop. This result suggests that CLIP’s effectiveness stems primarily from its generalisation ability rather than memorisation. Accordingly, our experiments emphasise this generalisation property, ensuring that the use of CLIP within our framework remains valid.
@@ B.3.2 Evidence for tabular datasets
For tabular datasets, we use the output of two classical unsupervised AD methods as evidence functions $T(x)$, namely, Local Outlier Factor (LOF) (Breunig et al.,, 2000) and Isolation Forest (IForest) (Liu et al.,, 2012).
@@
$T(x)=\frac{\sum_{x_{i}\in N_{k}(x)}\frac{\text{LRD}_{k}(x_{i})}{\text{LRD}_{k}(x)}}{|N_{k}(x)|}.$
@@ C.2 Experiments on industrial use case
CSP plant dataset. For the industrial setting, we utilise the simulated dataset introduced by Patra et al., (2024), which is generated by training a variational autoencoder on real-world data collected from an operational CSP plant. The dataset consists of thermal images of solar panels captured using infrared (IR) cameras, distinguishing it from the semantic and sensory anomaly datasets, as the images lack semantic structure and do not depict specific objects.
@@
Rule-based evidence. Foundation models, such as CLIP, which were previously used in our experiments on image datasets, are not applicable in specialised applications, such as detecting anomalous behaviour in solar power plants, due to the lack of semantic content in thermal images. This makes zero-shot methods like WinCLIP and AnoCLIP inapplicable. In contrast, while EPHAD can incorporate evidence from foundation models like CLIP, it also allows the seamless integration of domain-specific knowledge. To compute evidence, we utilise two of the four rules proposed by Patra et al., (2024) that indicate normal operational behaviour of the CSP plant. The first rule (R1) is based on the difference between consecutive images. Under normal conditions, the plant’s temperature is expected to remain relatively stable; therefore, substantial deviations from one image to the next suggest potential anomalies. To quantify this, pixel-wise squared differences are computed between every pair of consecutive images, and the 95th percentile of these differences is extracted as the representative evidence for each pair. The second rule (R2) involves the difference from the average daily temperature. Here, samples with average temperatures significantly diverging from the typical daily average could indicate anomalous behaviour. For this, the mean temperature of each day is first determined, and then the absolute difference between each image’s average temperature and that day’s mean is computed to serve as the evidence.
@@
Results. The results presented in Table 8 underscore the effectiveness and adaptability of our approach. Under a $10\%$ contamination setting, the baseline method ForecastAD experiences a performance drop of approximately $5\%$. However, by incorporating domain-specific rules R1 and R2 as sources of evidence using EPHAD and further using EPHAD-Ada, the performance nearly matches that on the clean dataset. It emphasises the value of leveraging structured, context-aware evidence to enhance the detection of anomalies. Importantly, foundation models like CLIP are unsuitable in this context due to the lack of semantic content in thermal imagery, rendering zero-shot approaches such as WinCLIP (Jeong et al.,, 2023) and AnoCLIP (Zhou et al.,, 2024) ineffective. EPHAD addresses this limitation by providing a flexible framework that integrates both powerful foundation models, where applicable, and domain-specific knowledge when necessary. This versatility enables EPHAD to deliver robust performance across diverse real-world anomaly detection tasks while maintaining efficiency and ease of deployment.
@@ C.3 Comparison against LOE and SoftPatch
To ensure a comprehensive evaluation, we compare the performance of our proposed post-hoc framework against SoftPatch (Jiang et al.,, 2022) and both variants of LOE (Qiu et al.,, 2022). However, it is important to note that, unlike our approach, both SoftPatch and LOE modify the training process to account for contamination, making it inapplicable to pre-trained networks without access to the training dataset and pipeline, which is our main focus.
@@
First, for comparison with LOE, we conduct experiments using the Neural Transformation Learning-based (NTL) AD method (Qiu et al.,, 2021) and evaluate it under four configurations: “Blind”, “Refine”, LOE-Hard and LOE-Soft. Additionally, we follow the same setup as LOE by extracting image features using pre-trained ResNet152 and WideResNet50 for semantic and sensory datasets, respectively, which are then used to train NTL. The results, summarised in Table 9, show that given a good evidence function, i.e. the performance of the evidence is better than the “Blind” configuration, our simple test-time framework outperforms LOE. Results on MVTec, CIFAR10, FMIST, and SVHN are examples of this behaviour. Also, on the ViSA dataset, the performance improves over the “Blind” and “Refine” configurations. In the converse situations where the performance of the evidence is lower than the “Blind” configuration, we observe a reduction in performance which can be accounted for by putting more emphasis on the AD model by adjusting $\beta$.
Now, we compare it against SoftPatch, an approach built upon PatchCore (Roth et al.,, 2022). SoftPatch enhances PatchCore by incorporating traditional anomaly detection (AD) techniques to refine the memory bank, specifically by identifying and re-weighting patches based on their outlier scores during training. While this strategy improves performance, it introduces a strong dependency on the choice of AD method and increases the computational burden of the training pipeline.

### v2

@@ B.3.1 Evidence for visual datasets
For the evidence function in image-based AD, we propose using Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021), a robust large-scale framework that learns joint vision-language representations from web-collected image-text pairs. While CLIP has been explored in prior work as a zero-shot AD method (Jeong et al., 2023; Zhou et al., 2024), its performance varies across different datasets. Although CLIP excels in detecting anomalies in real-world image datasets such as CIFAR10, it faces significant challenges when applied to domain-specific datasets, particularly those used for industrial inspection, like MVTec. This limitation stems from the lack of domain-specific knowledge in CLIP’s pre-training. In this section, we describe how CLIP is integrated into EPHAD as an evidence function $T(x)$, leveraging its strengths while mitigating its limitations in specialized domains.
Given a dataset $\mathcal{D}:=\{(x_{j},t_{j})\}_{j=1}^{n}$, CLIP trains an image encoder $e_{i}$ and a text encoder $e_{t}$ using contrastive learning (Chen et al., 2020), maximizing the cosine similarity between $e_{i}(x_{j})$ and $e_{t}(t_{j})$ for all $(x_{j},t_{j})\in\mathcal{D}$. For an input image $x$, CLIP performs zero-shot classification (Radford et al., 2021) by computing a $k$-way categorical distribution over a set of candidate class texts ${\mathcal{C}}=\{c_{1},\dots,c_{k}\}$
@@
Building on Jeong et al. (2023), we use CLIP as evidence function $T(x)$ in EPHAD. We start by defining two lists of textual prompt templates, ${\mathcal{T}}_{N}=\{n_{1},\cdots,n_{k}\}$ and ${\mathcal{T}}_{A}=\{a_{1},\cdots,a_{k}\}$, corresponding to normal and anomalous classes, respectively. The list of prompts is provided in Table 3. These templates are dataset-dependent, reflecting subjectivity (e.g., “missing wire" as anomalous for cables). For each label, we generate two lists of prompts for normal and anomalous cases using ${\mathcal{T}}_{N}$ and ${\mathcal{T}}_{A}$ and compute the mean of text embeddings $t_{N}$ and $t_{A}$. Finally, given an input image $x$, the evidence $T(x)$ during test-time is computed as:
$T(x):=\frac{\exp\left(\langle e_{i}(x),t_{N}\rangle/\gamma\right)}{\exp\left(\langle e_{i}(x),t_{N}\rangle/\gamma\right)+\exp\left(\langle e_{i}(x),t_{A}\rangle/\gamma\right)}.$
One potential concern when using pre-trained models like CLIP is the overlap between their training data and the test samples encountered in downstream tasks. Such overlap could challenge the assumption that test-time statistics are based solely on test data. However, Radford et al. (2021) provides an extensive analysis of this issue and shows that excluding all overlapping samples from CLIP’s pre-training corpus leads to only a negligible performance drop. This result suggests that CLIP’s effectiveness stems primarily from its generalisation ability rather than memorisation. Accordingly, our experiments emphasise this generalisation property, ensuring that the use of CLIP within our framework remains valid.
@@ B.3.2 Evidence for tabular datasets
For tabular datasets, we use the output of two classical unsupervised AD methods as evidence functions $T(x)$, namely, Local Outlier Factor (LOF) (Breunig et al., 2000) and Isolation Forest (IForest) (Liu et al., 2012).
@@
$T(x)=-\frac{\sum_{x_{i}\in N_{k}(x)}\frac{\text{LRD}_{k}(x_{i})}{\text{LRD}_{k}(x)}}{|N_{k}(x)|}.$
@@ C.2 Experiments on industrial use case
CSP plant dataset. For the industrial setting, we utilise the simulated dataset introduced by Patra et al. (2024), which is generated by training a variational autoencoder on real-world data collected from an operational CSP plant. The dataset consists of thermal images of solar panels captured using infrared (IR) cameras, distinguishing it from the semantic and sensory anomaly datasets, as the images lack semantic structure and do not depict specific objects.
@@
Rule-based evidence. Foundation models, such as CLIP, which were previously used in our experiments on image datasets, are not applicable in specialised applications, such as detecting anomalous behaviour in solar power plants, due to the lack of semantic content in thermal images. This makes zero-shot methods like WinCLIP and AnoCLIP inapplicable. In contrast, while EPHAD can incorporate evidence from foundation models like CLIP, it also allows the seamless integration of domain-specific knowledge. To compute evidence, we utilise two of the four rules proposed by Patra et al. (2024) that indicate normal operational behaviour of the CSP plant. The first rule (R1) is based on the difference between consecutive images. Under normal conditions, the plant’s temperature is expected to remain relatively stable; therefore, substantial deviations from one image to the next suggest potential anomalies. To quantify this, pixel-wise squared differences are computed between every pair of consecutive images, and the 95th percentile of these differences is extracted as the representative evidence for each pair. The second rule (R2) involves the difference from the average daily temperature. Here, samples with average temperatures significantly diverging from the typical daily average could indicate anomalous behaviour. For this, the mean temperature of each day is first determined, and then the absolute difference between each image’s average temperature and that day’s mean is computed to serve as the evidence.
@@
Results. The results presented in Table 8 underscore the effectiveness and adaptability of our approach. Under a $10\%$ contamination setting, the baseline method ForecastAD experiences a performance drop of approximately $5\%$. However, by incorporating domain-specific rules R1 and R2 as sources of evidence using EPHAD and further using EPHAD-Ada, the performance nearly matches that on the clean dataset. It emphasises the value of leveraging structured, context-aware evidence to enhance the detection of anomalies. Importantly, foundation models like CLIP are unsuitable in this context due to the lack of semantic content in thermal imagery, rendering zero-shot approaches such as WinCLIP (Jeong et al., 2023) and AnoCLIP (Zhou et al., 2024) ineffective. EPHAD addresses this limitation by providing a flexible framework that integrates both powerful foundation models, where applicable, and domain-specific knowledge when necessary. This versatility enables EPHAD to deliver robust performance across diverse real-world anomaly detection tasks while maintaining efficiency and ease of deployment.
@@ C.3 Comparison against LOE and SoftPatch
To ensure a comprehensive evaluation, we compare the performance of our proposed post-hoc framework against SoftPatch (Jiang et al., 2022) and both variants of LOE (Qiu et al., 2022). However, it is important to note that, unlike our approach, both SoftPatch and LOE modify the training process to account for contamination, making it inapplicable to pre-trained networks without access to the training dataset and pipeline, which is our main focus.
@@
First, for comparison with LOE, we conduct experiments using the Neural Transformation Learning-based (NTL) AD method (Qiu et al., 2021) and evaluate it under four configurations: “Blind”, “Refine”, LOE-Hard and LOE-Soft. Additionally, we follow the same setup as LOE by extracting image features using pre-trained ResNet152 and WideResNet50 for semantic and sensory datasets, respectively, which are then used to train NTL. The results, summarised in Table 9, show that given a good evidence function, i.e. the performance of the evidence is better than the “Blind” configuration, our simple test-time framework outperforms LOE. Results on MVTec, CIFAR10, FMIST, and SVHN are examples of this behaviour. Also, on the ViSA dataset, the performance improves over the “Blind” and “Refine” configurations. In the converse situations where the performance of the evidence is lower than the “Blind” configuration, we observe a reduction in performance which can be accounted for by putting more emphasis on the AD model by adjusting $\beta$.
Now, we compare it against SoftPatch, an approach built upon PatchCore (Roth et al., 2022). SoftPatch enhances PatchCore by incorporating traditional anomaly detection (AD) techniques to refine the memory bank, specifically by identifying and re-weighting patches based on their outlier scores during training. While this strategy improves performance, it introduces a strong dependency on the choice of AD method and increases the computational burden of the training pipeline.